OpenStudy (anonymous):

A right triangle with a median. The hypotenuse is y and where the median intersects the original right triangle has a segment 2 the whole hypotenuse is x. The short leg of the right triangle is 6. Find the value of the variables

OpenStudy (anonymous):

I'm not sure how to put a picture on here

OpenStudy (anonymous):

something similar (if anyone can help me) y + 2/6 = 6/2

OpenStudy (anonymous):

2y + 4 - 36 = 0 I'm not sure how to solve this???

OpenStudy (anonymous):

2(y+2- 18) = 0

OpenStudy (anonymous):

Can anyone help me with this please?